OpenStudy (anonymous):

hey me again answer this quickly and of u do ill give u medals promise no scam Which of the following rational #'s are equal to 12/20, -3/-5, 0.3 or 6/10 and question number two Which of the following rational #'s are equal to 12/15 4/5, 40/50, - 8/10 or 8/10

jimthompson5910 (jim_thompson5910):

12/20 = -3/-5 12/20 = 6/10 --------------------------- 12/15 = 4/5 12/15 = 40/50 12/15 = 8/10

OpenStudy (curry):

4/5 40/50 8/10

OpenStudy (anonymous):

r u sure these r it curry ill give u medal even if u take ur time

OpenStudy (anonymous):

oh and what about 3/5 the options r : 12/30, -3/-5, 0.3 or 6/10

OpenStudy (anonymous):

u 2 jim

OpenStudy (anonymous):

ill give u a medal

jimthompson5910 (jim_thompson5910):

3/5 = -3/-5 3/5 = 6/10 3/5 = 12/20 ... assuming you made a typo in the directions

OpenStudy (anonymous):

r these the right answer or not

jimthompson5910 (jim_thompson5910):

yes, they are the correct answers.

OpenStudy (anonymous):

oh ok ill be a fan

OpenStudy (anonymous):

thxs i got 2 go bye people